OpenStudy (keana):

Which expression is the simplified form of √a^13? a: a^5√a^3 b: a^12√a c: a^3√a^4 d: a^6√a

OpenStudy (anonymous):

\[\sqrt{a^{13}} = \sqrt{a^{2}\times a^{2}\times a^{2}\times a^{2}\times a^{2}\times a^{2}\times a}\]Take out all of the a^2 and leave them as a outside the radical.

OpenStudy (anonymous):